find the equation
slope= 3, passes through (1,2)

Answer by Alan3354(69443)

find the equation
slope= 3, passes through (1,2)
---------------
y-y1 = m*(x-x1) where (x1,y1) is (1,2) and m is the slope
y-2 = 3*(x-1)
y-2 = 3x-3
y = 3x - 1 (slope-intercept form)
3x - y = 1 (standard form)